Hello,
I'm working on an assignment for my Stats class. I am stumped on how to calculate this question. I believe it has to do with Binomial Distribution but I'm not 100% positive.
In a market study for Zellers, a researcher found that 70% of customers are repeat customers. If 12 customers are selected at random, find the probability of getting.
a. Exactly 9 of them are repeat customers
b. At least 9 of them are repeat customers
c. At most 9 of them are repeat customers
Thank you.

Answer by ewatrrr(24785)

 
Hi,
Binomial Distribution; Yes (repeat or not repeat)
p = .70, q = .30 and n = 12
P(x = 9) = 12C9(.70)^9(.30)^3 binompdf(12,.70,9)
P(x ≥ 9 ) = P(x=9) + P(x=10) + P(x=11) + P(x=12) 1 - binomcdf(12,.70,8)
P(x ≤ 9) = P(x=0) + P(x=1) + P(x=2).... + P(x=9) binomcdf(12,.70.9)